state and prove the necessary conditions for the existence of maximum value and minimum value?


1
Expert's answer
2021-09-17T00:28:05-0400

The conditions for a function to be maximum or minimum include if

"f\u2019 (x) = 0";

and for maximum "f\u2019\u2019 (x)" = negative

and for minimum "f\u2019\u2019 (x)" = positive.



Assume that "f: (a,b) \to \R"


Theorem:

Let "c \u2208 (a, b)" and f be continuous at c. If for some "\u03b4 > 0," f has a minimum increasing on "(c \u2212 \u03b4, c)" and decreasing on "(c,c+\u03b4)," then f has a local maximum at c.


Proof:

Choose any "x_1" and "x" such that c "\u2212 \u03b4 < x1 < x < c." Then "f(x_1) \u2264 f(x)" and by the continuity of "f" at "c" we have "f(x_1) ≤ \lim

x→c−

f(x) = f(c)."


Similarly, if "c < x_2 < c + \u03b4" then "f(x_2) ≤ \lim x→c+ f(x) = f(c)." This proves the result.